# Content-based Wake-up for Energy-efficient and Timely Top-$k$ IoT Sensing Data Retrieval

### v1

@@ V-C2 Minimum Energy Consumption
In order to compare the performance of CoWu and RR-scheduling, in terms of minimum energy consumption under the $k$-QAoI constraint, we obtain the optimal parameter set of $\{V_{\mathrm{th}},\zeta\}$ in Eq. (30), where $\gamma_{\mathrm{th}}$ is set to the $k$-QAoI of RR-scheduling, $\bar{\Delta}_{k}^{\mathrm{RR}}(N)$. Fig. 9 shows the minimum total energy consumption of CoWu obtained by Eq. (30) with RR-scheduling against $N$, when they are applied for timely top-5 data collection, where we set $e_{c}=0$ and $\Gamma$ = 1000, 5000 for the linear age case, in Eq. (4). In Fig. 9, total energy consumption of CoWu for the number of nodes equal to or less than 40 are not plotted. This is because CoWu can not realize smaller $k$-QAoI than RR-scheduling over this range of number of nodes. However, we can see that as the number of nodes becomes larger, CoWu can satisfy the constraint of $k$-QAoI, in which minimum total energy consumption of CoWu becomes smaller as $N$ increases. This is because as the number of nodes increases, the $k$-QAoI of round robin scheduling becomes larger, by which the sink can select the appropriate parameter $\zeta$ with a relatively larger threshold of $V_{\mathrm{th}}$ for CoWu signaling under the constraint of $k$-QAoI of RR-scheduling. Further, from Fig. 9, we can also see that the minimum total energy consumption of CoWu becomes larger as $\Gamma$ increases. This is because, as $\Gamma$ increases, the sink is required to set a relatively smaller threshold of $V_{\mathrm{th}}$ in order to avoid the increase of $k$-QAoI by the penalty of $\Gamma$, by which total energy consumption of sensor nodes increases. Finally, we can see that CoWu can achieve smaller energy consumption for larger number of nodes while satisfying the constraint of $k$-QAoI, thanks to the suppression of wake-up of nodes that observe unrelated data for the top-$k$ query. From this result, we can clearly see the effectiveness of CoWu against RR-scheduling in terms of $k$-QAoI and total energy consumption.
[…]
@@ V-C3 Maximum $k$
Fig. 10 shows the results of maximum $k$ against the different parameter settings. First, we can see that the value of $k$ becomes larger as the number of nodes increases. In RR-scheduling, data staleness for the nodes that transmit data very early time can not be avoided to realize reliable data transmission, which increases the average $k$-QAoI and its negative effect becomes larger as the number of nodes increases. On the other hand, CoWu achieves both high information freshness and high energy efficiency by carefully tuning the parameters $\{\zeta$, $V_{\mathrm{th}}\}$, especially when the value of $k$ is small.
@@
Furthermore, we can see that the maximum $k$ of setting 2 is equal to or larger than that of setting 1. When the single packet error occurs, $k$-QAoI of RR-scheduling becomes larger by the penalty of $\Gamma$, as it decreases the reliability of data collection. That is why the maximum $k$ of setting 2 becomes larger than that of setting 1 when the number of nodes is 20.
@@
From these observations, we can conclude that CoWu outperforms RR-scheduling in the context of timely top-$k$ data collection in the IIoT scenario in terms of total energy consumption and $k$-QAoI when the number of nodes equal to or larger than 60. We also calculate the ratio of the maximal value of $k$ against $N$, which we call the maximum $k$-$N$ ratio as in [5], based on the results obtained in Fig. 10, which shows the maximum region that the CoWu outperforms the RR-scheduling in terms of $k$-QAoI and total energy consumption. With our calculation, we confirmed that the maximum $k$-$N$ ratio reaches 0.2.
2020202040404040606060608080808010010010010000101010102020202030303030Number of Nodes $N$max $k$Setting 1: ($\Gamma,e_{\text{c}}$) = (1000, 0)Setting 2: ($\Gamma,e_{\text{c}}$) = (1000, 0.1)Setting 3: ($\Gamma,e_{\text{c}}$) = (5000, 0)Setting 4: ($\Gamma,e_{\text{c}},\alpha$) = (1000, 0, 0.02)
Figure 10: Maximum $k$ for CoWu against number of nodes with a different parameters settings.
@@ VI Conclusions
Possible avenues for future work include designing a wake-up control introducing the inference mechanism of the physical process to decide the optimal timing of sampling, wake-up, and data transmission for sensor nodes. Designing the wake-up control aiming the improvement of the model accuracy on the physical process is also an interesting direction.

### v2

@@ V-C2 Minimum Energy Consumption
In order to compare the performance of CoWu and RR-scheduling, in terms of minimum energy consumption under the $k$-QAoI constraint, we obtain the optimal parameter set of $\{V_{\mathrm{th}},\zeta\}$ in Eq. (33), where $\gamma_{\mathrm{th}}$ is set to the $k$-QAoI of RR-scheduling, $\bar{\Delta}_{k}^{\mathrm{RR}}(N)$. Fig. 10 shows the minimum total energy consumption of CoWu obtained by Eq. (33) with RR-scheduling against $N$, where we set $e_{c}=0$, $k=5$ and $\Gamma$ = 1000, 5000 for the linear age case in Eq. (4). In Fig. 10, total energy consumption of CoWu for the number of nodes or less than 40 is not plotted. This is because CoWu can not realize smaller $k$-QAoI than RR-scheduling over this range of number of nodes. However, we can see that as the number of nodes becomes larger, CoWu can satisfy the constraint of $k$-QAoI, in which minimum total energy consumption of CoWu becomes smaller as $N$ increases. This is because as the number of nodes increases, the $k$-QAoI of RR-scheduling becomes larger, by which the sink can select the appropriate parameter $\zeta$ with a relatively larger threshold of $V_{\mathrm{th}}$ for CoWu signaling under the constraint of $k$-QAoI of RR-scheduling. Further, from Fig. 10, we can also see that the minimum total energy consumption of CoWu becomes larger as $\Gamma$ increases. This is because, as $\Gamma$ increases, the sink is required to set a relatively smaller threshold of $V_{\mathrm{th}}$ in order to avoid the increase of $k$-QAoI by the penalty of $\Gamma$, by which total energy consumption of sensor nodes increases. Finally, we can see that CoWu can achieve smaller energy consumption for larger number of nodes while satisfying the constraint of $k$-QAoI, thanks to the suppression of wake-up of nodes that observe unrelated data for the top-$k$ query. From this result, we can clearly see the effectiveness of CoWu against RR-scheduling in terms of $k$-QAoI and total energy consumption.
00202020204040404060606060808080801001001001000020202020404040406060606080808080100100100100Number of Nodes $N$Minimum energy [mJ]CoWu (Theory) $\Gamma=1,000$CoWu (Theory) $\Gamma=5,000$RR-scheduling (Theory)
@@ V-C3 Maximum $k$
Fig. 11 shows the results of maximum $k$ against the different parameter settings, where we adopt the optimal transmission probability $p_{\mathrm{opt}}^{*}$ obtained in Eq. (32). First, we can see that the value of $k$ becomes larger as the number of nodes increases. In RR-scheduling, data staleness for the nodes that transmit data very early time can not be avoided to realize reliable data transmission, which increases the average $k$-QAoI and its negative effect becomes larger as the number of nodes increases. On the other hand, CoWu achieves both high information freshness and high energy efficiency by carefully tuning the parameters $\{\zeta$, $V_{\mathrm{th}}\}$, especially when the value of $k$ is small.
@@
Furthermore, we can see that the maximum $k$ of setting 2 is equal to or lower than that of setting 1 when $N>20$. In CoWu, a longer contention period is required when a single packet error happens, which increases the $k$-QAoI and total energy consumption. That is why the maximum $k$ of setting 2 is equal to or lower than that of setting 1. On the other hand, when the number of nodes is 20, maximum $k$ of setting 2 becomes larger than that of setting 1. This is because when the single packet error occurs, $k$-QAoI of RR-scheduling becomes larger by the penalty of $\Gamma$, as it decreases the reliability of data collection.
@@
From these observations, we can conclude that CoWu outperforms RR-scheduling in the context of timely top-$k$ data collection in the IIoT scenario in terms of total energy consumption and $k$-QAoI when the number of nodes equal to or larger than 60. We also calculate the ratio of the maximal value of $k$ against $N$, which we call the maximum $k$-$N$ ratio as in [5], based on the results obtained in Fig. 11, which shows the maximum region that the CoWu outperforms the RR-scheduling in terms of $k$-QAoI and total energy consumption. With our calculation, we confirmed that the maximum $k$-$N$ ratio reaches 0.25.
20202020404040406060606080808080100100100100002020202040404040Number of Nodes $N$max $k$Setting 1: ($\Gamma,e_{\text{c}}$) $=$ (1000, 0)Setting 2: ($\Gamma,e_{\text{c}}$) $=$ (1000, 0.1)Setting 3: ($\Gamma,e_{\text{c}}$) $=$ (5000, 0)Setting 4: ($\Gamma,e_{\text{c}},\alpha$) $=$ (1000, 0, 0.02)
Figure 11: Maximum $k$ for CoWu against the number of nodes with different parameters settings with $p_{\mathrm{opt}}^{*}$ in Eq. (32).
@@ VI Conclusions
Possible avenues for future work include designing a wake-up control introducing the inference mechanism of the physical process to decide the optimal timing of sampling, wake-up, and data transmission for sensor nodes. Designing the wake-up control aiming for the improvement of the model accuracy on the physical process is also an interesting direction.